A body of mass 10 kg is to be given an acceleration. A body of mass 10 kg is to be given an acceleration of 30 ms-2. Calculate the force required when the acceleration is

i. vertically upwards,

ii. horizontal and the body moves over a surface whose coefficient is 0.5
1
Expert's answer
2020-06-25T09:35:48-0400

i)


"F=m(a+g)=10(30+9.8)=398\ N"

ii)

"F=m(a+\mu g)=10(30+(0.5)9.8)=349
\ N"